Find the standard form of the hyperbola with verticies (0,3) and (0,-3) and asymptotes y=3/2x and y=-3/2x
Answer by lwsshak3(11628)

Find the standard form of the hyperbola with verticies (0,3) and (0,-3) and asymptotes y=3/2x and y=-3/2x
**
Given hyperbola has a vertical transverse axis:
Equation: (y-k)^2/a^2-(x-h)^2/b^2=1, (h,k)=(x,y) coordinates of center.
For given hyperbola:
center: (0,0)
length of vertical transverse axis=6 (-3 to 3)=2a
a=3
a^2=9
..
slope of asymptotes for hyperbolas with vertical transverse axis=a/b=3/2
b=a/3/2=2
b^2=4
..
Equation of given hyperbola:
y^2/9-x^2/4=1
see graph below as a good visual check of what the curve should look like
y=±(9x^2/4+9)^.5